Day 04-Jul-1999 Shift Report

Mark Timko (timko@ktev3)
Sun, 4 Jul 1999 15:30:48 -0500 (CDT)

Shift Summary for Day shift, 04/Jul/99

On Shift: Arun Tripathi, Mark Timko.

Hours of beam available at all intensities: 05 hr, 40 min,
Hours of beam available at nominal intensity: 05 hr, 38 min,
Hours of E832 data (special runs not included): 05 hr, 35 min,
Hours of special runs (laser scans, muon runs...):00 hr, 50 min.

Accounting of lost time:
Tapes full; ended run (~ 5 minutes of beam lost)
Linac RF station water leak. Beam down for 2 hours and 50 minutes.

========================================================================
Runs ended this shift:
Run 13812 - 10.4M events - BEAM_832
Run 13814 - no data beam down.
Run 13815 - - LASER_SCAN

Tapes brought to FCC this shift:
QKB029, QKC083, QKC084, QKC085, QKK083, QKK084, QKK085,
QKS083, QKS084, QKS085,

========================================================================
Accesses during shift: None (although we wanted to)

========================================================================
Overview of shift:
E832 running, no beam for ~ 3 hours but the MCR kept estimating the
fix at 30 minutes so this precluded us from making an access.
Did a Long laser scan.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13812 09:57 Onsplit: Mounted tape QKN007 on KTEV4, drive tps3d2.

13812 197 10:07 Low SEM = 2.4E12. Main Control Room says
they are timing the beam for losses and that the
rate should be back up in ~ 5 minutes.

13812 200 10:10 SEM = 5E12.

13812 208 10:23 DYC 39 (laser) busy. Couple of KFRENDs fixed it.

13812 10:25 Tapes fill up. End run.

in between 10:30 Linac RF station 1 has a PA crowbar (but we don't
know this yet).

13814 10:44 New Run, new tapes QKB030, QK*086 to QK*088.
But no beam! Called MCR and they say about 15 minutes.

13814 11:09 MCR says Linac RF station water leak; expert enroute.
Estimate is 30 minutes.

13814 11:20 New estimate is 45 minutes. So we stopped the run to
take a laser scan run.

13815 11:25 Long laser scan run. Done at 12:13

13816 12:20 Start new run. Beam still down.

13816 43 13:20 Beam Returns.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

1. Drive 3 on ktev2 is broken. Drive 5 is being used instead. Keep this in
mind while mounting mew tapes.

2. Ed B. says don't do a muon run tonight; wait until tomorrow.

[End of Shift Report]